please help me solve, [4x-1]-[2x+2]=x+5
Answer by dave41082(4)

Begin by eliminating the brackets using the distributive property:
4x - 1 - 2x - 2 = x + 5
Then subtract x from both sides to get:
4x - 1 - 2x - 2 - x = 5
Then, combine like terms on the left side:
x - 3 = 5
Then add 3 to both sides to get:
x = 8